# Comparing growth of titania and carbonaceous dusty nanoparticles in weakly magnetised capacitively coupled plasmas

### v1

@@ 3.2 Laser light scattering
Figure 7: Evolution of the dust cloud of Ar/TTIP, without magnetic field $\left(T_{c}\approx 77s\right)$. Dust cloud at (a) $T_{c}/4$, (b) $T_{c}/2$, (c) $43T_{c}/$ 4, and (d) $T_{c}$. (e) Squares, where light emission intensity was calculated, overlaid on (b) spanning the dust cloud. (f) Cyclic behavior of light emission from 3 growth cycles of the blue square. (g) NFT of the cycles in the blue square showing a frequency $(1/T_{c})$ of 0.0130 Hz, which is also the most occurring frequency, i.e. the mode from all the 30 squares. (h) Intensity of the mode frequency measured from all the 30 squares and spanning the dust cloud overlaid on (b).
Figure 8: Evolution of the dust cloud of Ar/$C_{2}H_{2}$, without magnetic field. The mode frequency is 0.0087 Hz and $T_{c}\approx 115s$. All the subfigures follow the same description and data analysis similar to Fig. 7.
Figure 9: Evolution of the dust cloud of Ar/TTIP, with magnetic field. The mode frequency is 0.0315 Hz and $T_{c}\approx 32s$. All the subfigures follow the same description and data analysis similar to Fig. 7.
Figure 10: Evolution of the dust cloud of Ar/$C_{2}H_{2}$, with magnetic field. The mode frequency is 0.0264 Hz and $T_{c}\approx 38s$. All the subfigures follow the same description and data analysis similar to Fig. 7.
@@
The laser light that was scattered from the dust cloud was recorded using a CMOS camera. The time frame of the CMOS camera was set from the frame rate, whereby a frame was recorded every 20 ms, i.e 50 FPS. The first frame with plasma on was set to t = 0. In order to compare the dust cloud to each other at the same time during their cycles, 4 images are chosen and shown. They show the region the region of the dust cloud between the electrodes at $T_{c}/4$, $T_{c}/2$, $43T_{c}/$ 4, and $T_{c}$, where $T_{c}$ was the cycle time according to the OES. These are shown in Fig. 7 - 10 (a), (b), (c), and (d). Figs. 7 and 8 present the evolution of the Ar/TTIP and Ar/$C_{2}H_{2}$ dust clouds in the absence of a magnetic field, while Figs. 9 and 10 present their evolution with a magnetic field, respectively.
@@
30 boxes, 100 $\times$ 100 pixels each, were drawn on the dust cloud, 10 in 3 different rows. The boxes were drawn in order to cover the entire region of the cloud between the electrodes where light scattering from dust particles is observed. Within each analysis box, the average light intensity was calculated from all of the pixels in that region. A time series, spanning over three growth cycles, was then created by recording the light intensity as a function of time for each box. A numerical Fourier transform (NFT) of the time series was performed for each box to identify the dominant frequencies. The most occuring dominant frequency, i.e. mode frequency, measured from the 30 NFT was chosen to represent and calculate the growth cycles. In Fig. 7 - 10, the 30 boxes are shown in (e), examples of the cyclic variation of the light emission intensity are shown in (f), and examples of the mode frequency are shown in (g). The X-axis of (f) indicates the time calculated from the camera frames. Moreover, (h) shows the magnitude of the mode frequency in the 30 boxes superimposed on the dust cloud at $T_{c}/2$. For the Ar/TTIP dusty plasma, without the presence of the magnetic field, the mode frequency and cycle time were 0.0130 Hz and 77 s respectively, and with the presence of the magnetic field, the mode frequency and cycle time were 0.0315 Hz and 32 s respectively. For the Ar/$C_{2}H_{2}$ dusty plasma, without the presence of the magnetic field, the mode frequency and cycle time were 0.0087 Hz and 115 s respectively, and with the presence of the magnetic field, the mode frequency and cycle time were 0.0264 Hz and 38 s respectively.
@@ 3.3 Size Distribution
The particles reached their maximum size by the end of each cycle. When a new cycle started, new particles initiated their growth. This growth exhibited a consistent linear progression in size throughout two cycle, as depicted in Figure 11. The time frame for the size distribution analysis was determined by referring to the frames from CMOS camera images. Each particle growth experiment was meticulously recorded and timed according to the number of frames, spanning from the activation of plasma to its deactivation. Even though the titania growth cycle without the presence of the magnetic field ended after approximately 77 s, some dust from the previous cycle still remained in the plasma until about 110 s. The presence of the two cycles are shown in Fig. 1(c). However, this situation with the presence of two growth cycles is only observed in the Ar/TTIP dusty plasma without magnetic field. The maximum radius of the particles after 77 s were 235 $\pm$ 20 nm, however, the particles’ radii up grew to 355 $\pm$ 12 nm at 110 s, as shown in Fig. 11. The carbonaceous particles reached their maximum radius of 346 $\pm$ 14 after 110 s, shortly after which the new cycle started, as shown in Fig. 11. The maximum radii of titania and carbonaceous dust particles grown during the presence of magnetic field were 294 $\pm$ 22 nm and 200 $\pm$ 16 nm, respectively, as shown in Figs. 11 and 11.
@@ 4 Discussion
In prior investigations of dusty plasma particle growth, it was determined that a magnetic field of 1 Tesla reduced the length of time of the growth cycle [56]. Our findings are consistent with these earlier observations, even with a lower magnetic field, for both types of dust particles, as summarized in Table 1. Literature has shown that a weak magnetic field as low as 50 Gauss can affect the plasma potential, and hence the electric field in capacitively coupled Ar plasmas [66]. The smaller radii of nanoparticles grown during the presence of magnetic field suggest that there might be a reduction in the electric field. This is consistent with preliminary measurement of our electric field, however a full description the effect of magnetic field on the electric field is requires a separate study. Since the force balance levitating the dust cloud prior to the end of the cycle depends primarily on electric and gravitational forces, the reduced electric field allows the particles to move out of the plasma quicker, and hence leading to the faster growth cycle.
@@
In addition to the faster growth cycles when the magnetic field is present, there is also a noticeable impact within the spatial distribution of the dust cloud in both dusty plasma during the presence of the magnetic field. Specifically, the presence of the magnetic field resulted in a larger void, for the growth of carbonaceous particles as compared to the growth of the titania particles. Boxes are drawn in the central region of the plasma to show the void in Fig. 9 (b) and (c) for titania, and 10 (b) and (c) for carbon. To visualize this effect better, a line profile of the emitted light intensity versus x (mm) was examined in the four dusty plasmas over two growth cycles. The analysis covered the electrode’s diameter, specifically within the range of $-37.5\leq$ x (mm) $\leq 37.5$, positioned at the midplane between the two electrodes at z = 12.5 mm, as shown in Fig. 2(c). The variation in intensity along the line, 50 pixels thick, represents the variation in spatial dust cloud distribution. The intensity obtained were stacked vertically to show the temporal evolution of the dust cloud. Fig. 11(a) and 11(b) illustrate this for Ar/TTIP and Ar/$C_{2}H_{2}$ dusty plasmas without a magnetic field, while Fig. 11(d) and 11(d) show the same for Ar/TTIP and Ar/$C_{2}H_{2}$ dusty plasmas with a magnetic field, respectively. The intensity for each dusty was normalized to its highest number to enable a more direct comparison to be made between the particle growth of the two dust particle types. The temporal intensity profile appears to be similar for both kinds of dusty plasma without magnetic field and dissimilar during the presence of the magnetic field.
@@
Figure 12: Temporal evolution of the light emission intensity from a line at z = 12.5 mm across the diameter of the dust cloud. Without magnetic field: (a) Ar/TTIP and (b) Ar/$C_{2}H_{2}$. With magnetic field: (d) Ar/TTIP and (e) Ar/$C_{2}H_{2}$. Sum of intensity of both dusty plasma from light emitted intensity due to laser scattering across two growth cycles: (c) without magnetic field, (f) with magnetic field.
To represent the overall radial distribution of the dust particles, the light intensity from Figs. 9 (b) and (c) and 10 (b) and (c) was summed over all frames for each horizontal positions, x. This is then used to create two line plots shown in Figs. 11(c) and 11(f) which indicates the overall spatial distribution of the grown dust particles over two growth cycles without and with the magnetic field. These plots serve as a quantitative comparison of the two dust clouds to reveal any similarities and differences. In the absence of a magnetic field, the radial intensity variation of the two dusty plasmas exhibits similarities. However, with a magnetic field present, the Ar/TTIP dusty plasma shows peaks at the center and edges of the electrode, whereas the Ar/$C_{2}H_{2}$ dusty plasma exhibits a peak only at the edges and a reduced intensity, due to a void, in the middle of the plasma. This is consistent with the visual observation that there was more titiania dust in the middle of the plasma than there was carbonaceous dust, i.e there was a difference between the spatial distribution of the two dust clouds. To understand why the dust cloud reacts differently to the presence of a magnetic field, it is important to explore changes in plasma conditions resulting from the presence of the magnets. Therefore, our investigation focuses on how the magnetic field has influenced plasma conditions, which can ultimately affect the spatial distribution of the dust cloud. Using a Langmuir probe in an Ar plasma, we measured the plasma parameters of floating potential and electron temperature . The results of the measurements are shown in Fig. 13. There is a radial variation in floating potential as shown in Fig. 12(a). However, there does not seem to be a radial variation of electron temperature.
@@
We now consider the effects of the magnetic field on the dust. In this experiment, only the electrons were magnetised during the presence of the magnetic field, according to the Hall parameter, a commonly employed metric for calculating the degree magnetization of species in a dusty plasma due to their charge accumulation [67, 68]. Ions and dust particles are not magnetized. Therefore, there is no magnetization due to charge accumulation on the dust that can explain the differences in spatial dust distribution of titania and carbonaceous nanoparticles. According to Fig. 4(b), the magnetic field, $B_{z}$ is constant in the region - 15 $\leq$ x(mm) $\leq$ 15. However, there is a change in $B_{z}$ with respect to x outside that range until the edge of the electrodes. It is plausible that a gradient in the magnetic field is contributing to the differences seen between the spatial distribution of the two dusty plasma. Therefore, we hypothesize that the particles may be expressing alternative magnetic characteristics based on their material properties that might potentially account for variations in the spatial dust distribution and the subsequent dynamics of the dust clouds under the influence of the magnetic field. For example, numerous studies have shown that carbonaceous particles generated in the Ar/$C_{2}H_{2}$ plasma possess aromatic carbon bonds [56, 69, 70]. This is a property which is known to induce diamagnetism in carbonaceous materials [71, 72, 73]. It is possible that the diamagnetic property of the carbonaceous dust is repelling them away from the center of the plasma towards the edge, hence explaining the peak at the edge and a drop in intensity in the middle of the Ar/$C_{2}H_{2}$ dusty plasma. Additionally, it has been shown that the titania particles grown in Ar/TTIP plasma are amorphous [34]. Research have also shown that amorphous titania nanomaterials exhibit paramagnetic behaviors which originate from oxygen vacancies, surface adsorbents, and/or the disordered structure inherent in amorphous materials [74]. If the titania nanoparticles are exhibiting similar paramagnetic properties as in these studies, it is possible that they are being retained towards the middle of the plasma, hence contributing to the peak in intensity. However, in these low temperature plasmas, TTIP not only decomposes into $TiO_{2}$, but also into propene $(C_{3}H_{6})$ and isopropanol $(C_{3}H_{8}O)$, through a chemical pathway given by Eq. 1 [75].
$Ti\left(OC_{3}H_{7}\right)_{4}\rightarrow TiO_{2}+2C_{3}H_{6}+2C_{3}H_{8}O$ (1)
It is possible that some diamagnetic carbonaceous dust are grown from $C_{3}H_{6}$ and $C_{3}H_{8}O$. This may explain how the formation of dust might lead to a peak being seen at the edge of the Ar/TTIP dusty plasma, similar to the Ar/$C_{2}H_{2}$ dusty plasma. Future studies will explore the material properties of these dust particles and their correlation with magnetic field strength.
@@ 5 Conclusion
This paper compared the growth of titania and carbonaceous nanoparticles as dusty plasmas, in the presence and absence of a weak magnetic field. The growth in both dusty plasmas displays cyclic behavior, with a quicker cycle time when the magnetic field was present. The growth cycles were measured from OES of the plasma and NFT analysis of light emission intensity across the dust cloud. The particles reached their maximum size at the end of the cycle. The spatial distribution of the dust cloud looked similar for both kinds of dusty plasma without magnetic field. However, during the presence of the magnetic field, the two dust cloud responded differently. Measurement of background plasma parameters reveals changes in floating potential during the presence of the magnetic field. However, these measurements are not enough to explain the differences between the two dust cloud during the presence of the magnetic field. Although the dust are not magnetized according to the Hall parameter, carbonaceous dust and titania dust exhibited distinct responses to the applied magnetic field. Coincidentally, the magnetic field was relatively stronger at the middle of the plasma than at the edges, hence there was a gradient in the field. Therefore, these variations in the spatial distribution of the dust cloud may potentially be attributed to different magnetic properties of the dust to the gradient, such as paramagnetism in titania and diamagnetism in carbonaceous dust.

### v4

@@ 3.2 Laser light scattering
$\mathbf{T_{c}}$/4
$\mathbf{T_{c}}$/2
[…]
$\mathbf{T_{c}}$
Without Magnetic Field: Ar/TTIP
Without Magnetic Field: Ar/$C_{2}H_{2}$
With Magnetic Field: Ar/TTIP
With Magnetic
Field:
Ar/$C_{2}H_{2}$
Figure 7: Camera images of the four different dust cloud (row) at the same time (column) within each experiment, $T_{c}/4$, $T_{c}/2$, $43T_{c}/$ 4 and $T_{c}$, as labelled. The orange line is drawn at z = 12.5 mm of $T_{c}/2$, in order to compare the laser light scattering intensity in the mid-plane over 2 cycles in Fig. 10.
@@
The laser light scattering from the dust cloud was recorded using a CMOS camera. Dust density waves, which can be used to calculate dust density and temperature, were detected at the edge of the dusty plasma, however, they were not further studied herein [91]. The time frame was set from the frame rate, whereby a frame was recorded every 20 ms, i.e 50 FPS. The first frame with plasma on was set to t = 0. In order to compare the different dust cloud to each other at the same time during their cycles, 4 images are chosen and shown. The region of the dust cloud between the electrodes are shown in Fig. 7, with each column being at $T_{c}/4$, $T_{c}/2$, $43T_{c}/$ 4, and $T_{c}$, where $T_{c}$ was the cycle time according to the OES.
30 squares
Cycles from the blue square
Mode frequency
Magnitude of mode frequency
Without Magnetic Field: Ar/TTIP
Without Magnetic Field: Ar/$C_{2}H_{2}$
With Magnetic
Field:
Ar/TTIP
With Magnetic
Field:
Ar/$C_{2}H_{2}$
Figure 8: Each row represent an individual dusty plasma as labelled. The first column shows 30 squares of $\sim$ 100 $\times$ 100 pixels, in 3 rows and 10 columns, over the dust cloud at $T_{C}/2$. The second column shows the cyclic variation in light intensity at row 1 and column 3 (blue square). The third column shows the mode frequency obtained from all 30 square. The fourth column shows the magnitude of the mode frequency in all 30 squares overlayed over the dust cloud at $T_{C}/2$.
@@
Unlike OES which scans light emission across a volume, the laser light scattering occurs in a 2D plane. In order to consistently compare the cycle time, 30 squares measuring $\sim$ 100 $\times$ 100 pixels each were drawn on each dust cloud, 10 in 3 different rows as shown in the first column of Fig. 8. They were drawn to cover the entire region of the cloud between the electrodes where light scattering from dust particles is observed. Within each analysis box, the average light intensity was calculated from all of the pixels. A time series, spanning over three growth cycles, was then created by recording the light intensity as a function of time for each box. A numerical Fourier transform (NFT) of the time series was performed for each box to identify the dominant frequencies. The most occurring dominant frequency, i.e. mode frequency, measured from the 30 NFT was chosen to represent and calculate the growth cycles. Examples of the cyclic variation of the light emission intensity from the first row and third column (blue box) is shown in the second column of Fig. 8. The average intensity over 3 cycles is subtracted from the data before calculating NFT, hence the negative magnitudes. The X-axis of indicates the time calculated from the camera frames. The mode frequency from all 30 boxes is shown in the third column of Fig. 8. Moreover, the magnitude of the mode frequency in the 30 boxes superimposed on the dust cloud at $T_{c}/2$ is shown in the fourth column of Fig. 8. For the Ar/TTIP dusty plasma, without (with) the presence of the magnetic field, the mode frequency and cycle time were 0.0130 Hz (0.0315 Hz) and 77 s (32 s) respectively. For the Ar/$C_{2}H_{2}$ dusty plasma, without (with) the presence of the magnetic field, the mode frequency and cycle time were 0.0087 Hz (0.0264 Hz) and 115 s (38 s) respectively. The mode frequency and cycle time were rounded to two significant figures and they are in agreement with the values obtained from OES.
@@ 3.3 Size Distribution
In order to obtain the size of the nanoparticles, the individual experiments were ran for a certain instant of time and the nanoparticles subsequently were collected on a glass slide at the bottom electrode. The substrates were sputter coated with 15 nm of gold before being analyzed by Scanning Electron Microscope (SEM) (JEOL 7200F Field Emission SEM). The particles reached their maximum size by the end of each cycle. When a new cycle started, new particles initiated their growth. This growth exhibited a consistent linear progression in size throughout two cycle, as depicted in Figure 9. The time frame for the size distribution analysis was determined by referring to the frames from CMOS camera images. Each particle growth experiment was meticulously recorded and timed according to the number of frames, spanning from the activation of plasma to its deactivation.
At the end of the first cycle, the void expands, followed by a newer cycle which eventually replaces the precious one. Therefore, the size distribution of nanoparticles is reset in the second cycle. The growth of titania without magnetic field is the only exception, whereby, the dust from the previous cycle moves relatively slower out of the plasma and hence a bimodal size distribution is collected at 90 and 110 seconds. The maximum radius of the particles after 77 s were 235 $\pm$ 20 nm, however, the particles’ radii up grew to 355 $\pm$ 12 nm at 110 s, as shown in Fig. 9(a). The carbonaceous particles reached their maximum radius of 346 $\pm$ 14 after 110 s, shortly after which the new cycle started, as shown in Fig. 9(c). The maximum radii of titania and carbonaceous dust particles grown during the presence of magnetic field were 294 $\pm$ 22 nm and 200 $\pm$ 16 nm, respectively, as shown in Figs. 9(b) and 9(d).
@@ 4 Discussion
In prior investigations of dusty plasma particle growth, it was shown that a weak magnetic field of 320 Gauss reduced the growth cycle time [75]. Our findings are consistent with these observations for both types of dust particles, as summarized in Table 1. Literature has shown that a weak magnetic field as low as 50 Gauss can affect the plasma potential, and hence the electric field in capacitively coupled Ar plasmas [92]. The smaller radii of nanoparticles grown during the presence of magnetic field suggest that there might be a reduction in the electric field. This is consistent with preliminary measurement of our electric field, however a full description the effect of magnetic field on the electric field is requires a separate study. Towards the end of the growth cycle, the dominant forces acting on the nanoparticles are gravitational, electric and ion-drag. Therefore, a possible decrease in the electric force during the presence of the magnetic field is probably reducing the force balance, thus causing the faster growth cycle and reduced nanoparticles’ size.
@@
In addition to the faster growth cycles when the magnetic field is present, there is also a noticeable impact within the spatial distribution of the dust cloud in dusty plasma during the presence of the magnetic field. Specifically, the presence of the magnetic field resulted in a larger void, for the growth of carbonaceous particles as compared to the growth of the titania particles. Variations in chamber geometry are expected to influence the dust cloud. Here, there is a groove in the lower electrode to fit in the substrate for particle collection and a protrusion of the upper rf-electrode. However, this geometry is kept constant throughout all the experiments. Therefore, any changes in the dust cloud is not attributed to changes in geometry, but instead, to the presence of the magnetic field. Moreover, when the magnets are present, they are put either above the top electrode or below the bottom electrode, so that they do not change the geometry of the surfaces which are directly in contact with the plasma and do not interfere with the dust cloud.
@@
To visualize the overall radial distribution of the dust particles, the light intensity along the orange line of $T_{C}/2$ of Fig. 7 was summed over all frames for horizontal positions, in the range of $-37.5\leq$ x (mm) $\leq 37.5$, positioned at the midplane between the two electrodes at z = 12.5 mm, as shown in Fig. 3(c). The variation in intensity along the line, 50 pixels thick, represents the variation in spatial dust cloud distribution. The light intensity obtained was stacked vertically to show the temporal evolution of the dust cloud. Fig. 10(a) and 10(b) illustrate this for Ar/TTIP and Ar/$C_{2}H_{2}$ dusty plasmas without a magnetic field, while Fig. 10(d) and 10(d) show the same for Ar/TTIP and Ar/$C_{2}H_{2}$ dusty plasmas with a magnetic field, respectively. The intensity for each dusty was normalized to its highest number to enable a more direct comparison to be made between the particle growth of the two dust particle types. The temporal intensity profile appears to be similar for both kinds of dusty plasma without magnetic field and dissimilar during the presence of the magnetic field. Line intensity plots are shown in Figs. 10(c) and 10(f) which shows the overall spatial distribution of the grown dust particles over two growth cycles without and with the magnetic field. These plots serve as a quantitative comparison of the two dust clouds to reveal any similarities and differences. In the absence of a magnetic field, the radial intensity variation of the two dusty plasmas exhibits similarities. However, with a magnetic field present, the Ar/TTIP dusty plasma shows peaks at the centre and edges of the electrode, whereas the Ar/$C_{2}H_{2}$ dusty plasma exhibits a peak only at the edges and a reduced intensity, due to a void, in the middle of the plasma. This suggests that there was more titiania dust in the middle of the plasma than there was carbonaceous dust; thus, there was a difference between the spatial distribution of the two dust clouds at 500 Gauss.
@@
Figure 10: Temporal evolution of the light emission intensity from a line at z = 12.5 mm across the diameter of the dust cloud, as drawn in Fig. 7. Without magnetic field: (a) Ar/TTIP and (b) Ar/$C_{2}H_{2}$. With magnetic field: (d) Ar/TTIP and (e) Ar/$C_{2}H_{2}$. Sum of intensity of both dusty plasma from light emitted intensity due to laser scattering across two growth cycles: (c) without magnetic field, (f) with magnetic field.
To understand why the dust cloud reacts differently to the presence of a magnetic field, it is important to explore changes in plasma conditions resulting from the presence of the magnets. Therefore, our investigation focuses on how the magnetic field has influenced plasma conditions, which can ultimately affect the spatial distribution of the dust cloud. Using a Langmuir probe in the background Ar plasma, we measured the plasma parameters of floating potential and electron temperature. A voltage sweep was done between - 40 V and 40 V, from which the current was measured. The floating potential is defined as the voltage measured when the current collected by the probe tip is zero. By approximating a linear fit to the logarithm of the electron current versus voltage beyond the floating potential, the electron temperature was calculated from the inverse of the slope [93]. The results of the measurements are shown in Fig. 11. There is a radial variation in floating potential as shown in Fig. 11(a). However, there does not seem to be a radial variation of electron temperature. Even though these measurements were made in the absence of dust, we can suggest how the trend of the measurements will change. When the dust cloud has formed, the dust density is typically $\sim$ $10^{13}$ $m^{-3}$ [94]. This leads to an electron depletion in the background plasma due to the Havnes parameter being $>1$ [95]. We expect the overall floating potential of the plasma to slightly increase during the presence of the dusty nanoparticles[96]. At 500 Gauss, the changes in the floating potential, and consequently the radial electric field in the plasma, is possibly contributing to the faster cycle time. It is possible that the magnitude of several forces confining the particles in the plasma has decreased, thus leading to the faster cycle time. Additionally, it has already been shown that the electron temperature of particle growth as dusty plasma varies cyclically, in accordance with the dust growth cycles [78].
@@
We now consider the effects of the magnetic field on the dust. First, according to the Hall parameter, a commonly employed metric for calculating the degree magnetisation of charged species in a plasma [97, 98], only the electrons were magnetised at 500 Gauss. Ions and dust particles were not magnetised, due to their heavier masses. Therefore, there is no magnetisation due to charge accumulation on the dust that can explain the differences in spatial dust distribution of titania and carbonaceous nanoparticles. Second, there is a gradient in the radial direction of the magnetic field. According to Fig. 5(b), the magnetic field, $B_{z}$ is constant in the region - 15 $\leq$ x(mm) $\leq$ 15. However, there is a change in $B_{z}$ with respect to x outside that range until the edge of the electrodes. It is possible that the gradient is contributing to the differences seen between the spatial distribution of the two dusty plasma. A different study of the magnetic properties of the materials is necessary in the future in order to determine if they are contributing to the differences seen in the experiments due to the gradient.
@@ 5 Conclusion
We compared the growth of titania and carbonaceous nanoparticles as dusty plasmas, in the presence and absence of a weak magnetic field of 500 Gauss. Both particle growth displayed a cyclic behaviour, with a quicker cycle time at 500 Gauss. The cycle times were measured from OES of the plasma and NFT analysis of light emission intensity of the dust cloud. The particles reached their maximum size at the end of the cycle. The spatial distribution of the dust cloud appeared similar for both dusty plasma without magnetic field. However, during the presence of the magnetic field, the two dust cloud appeared differently. Measurement of background plasma parameters revealed changes in floating potential during the presence of the magnetic field. However, these measurements were not enough to explain the differences between the two dust cloud during the presence of the magnetic field. Furthermore, the ions and dust were not magnetized according to the Hall parameter. Nevertheless, there was a gradient in the radial distribution of magnetic field. Future studies need to investigate the material properties and whether the gradient is responsible for the differences in the spatial distribution of the dust cloud between titania and carbonaceous dusty plasmas.